A 45 kg load is being pulled vertically up by a rope whose breaking strength is 495N. With what maximum acceleration the load can be pulled up?

Grade:11

1 Answers

Arun
5 years ago
Dear student
 
T – mg = ma
 
495 – 45*10 = 45 * a(max)
 
a(max) = 1 m/sec^2
 
 
Regards
Arun (askIITians forum expert)